OpenStudy (anonymous):

Help please... What is the simplest form of the radical expression? sqrt2 + sqrt5 over sqrt2 - sqrt5

OpenStudy (driftracer305):

\[(\sqrt{2} + \sqrt{5}) (\sqrt{2} + \sqrt{5})/ (\sqrt{2}-\sqrt{5}) (\sqrt{2}+\sqrt{5})\]

OpenStudy (driftracer305):

\[(\sqrt{2}-\sqrt{5})(\sqrt{2}+\sqrt{5}) / 2-5\]

OpenStudy (driftracer305):

\[(2 + \sqrt{2}\sqrt{5}-\sqrt{5}\sqrt{2}-5) / -3\]

OpenStudy (driftracer305):

\[(-3 (\sqrt{2}\sqrt{5}-\sqrt{5}\sqrt{2})) -3\]